# SODAs: Sparse Optimization for the Discovery of Differential and Algebraic Equations

### v1

@@ Abstract
Differential-algebraic equations (DAEs) integrate ordinary differential equations (ODEs) with algebraic constraints, providing a fundamental framework for developing models of dynamical systems characterized by timescale separation, conservation laws, and physical constraints. While sparse optimization has revolutionized model development by allowing data-driven discovery of parsimonious models from a library of possible equations, existing approaches for dynamical systems assume DAEs can be reduced to ODEs by eliminating variables before model discovery. This assumption limits the applicability of such methods to DAE systems with unknown constraints and time scales. We introduce Sparse Optimization for Differential-Algebraic Systems (SODAs), a data-driven method for the identification of DAEs in their explicit form. By discovering the algebraic and dynamic components sequentially without prior identification of the algebraic variables, this approach leads to a sequence of convex optimization problems and has the advantage of discovering interpretable models that preserve the structure of the underlying physical system. To this end, SODAs improves numerical stability when handling high correlations between library terms—caused by near-perfect algebraic relationships—by iteratively refining the conditioning of the candidate library. We demonstrate the performance of our method on biological, mechanical, and electrical systems, showcasing its robustness to noise in both simulated time series and real-time experimental data.
@@ 1 Introduction
Data-driven discovery of dynamical models of complex systems promises to rapidly accelerate our understanding, from learning new equations governing cellular biomechanics [1] to the automatic generation of control models for robotics [2, 3]. A major challenge of many complex dynamical systems is the separation of timescales and the integration of physical constraints and conservation laws. Timescale separation is ubiquitous in complex biological and physical systems, from metabolic networks to multi-body robotics and power grids. Even in well-studied systems such as those following mass-action kinetics and deterministic Newtonian physics, numerical simulation and parameter estimation are challenging due to the inherent stiffness of such systems [4, 5, 6]. In this context, model selection is an even greater challenge as the structural form of the model, including its conservation laws and relevant timescales, may be unknown. In this work, we propose a method for data-driven discovery of complex systems by designing sparse-model selection within a differential-algebraic equation (DAE) framework, where conservation laws and quasi-steady-state equations are learned from time-series data as algebraic equations.
@@
Model identification or selection focuses on the general challenge of discovering the functional form of a model using data. For many dynamical systems—including those following mass-action kinetics or electromechanical conservation laws—it is reasonable to enumerate a library of possible interaction terms or candidate functions $\Theta=\{\theta_{1}(\mathbf{x}),\ldots,\theta_{J}(\mathbf{x})\}$, each representing a meaningful physical quantity as a function of the system state $\mathbf{x}(t)$. The key challenge is to downselect the most suitable subset of terms from this library. For example, a system of differential equations can often be described using a linear combination of nonlinear functions: $\dot{x}_{i}=\sum_{j=1}^{J}p_{j}^{i}\theta_{j}(\mathbf{x}),$ for each state variable $i=1,\ldots,d$, where $p_{i}^{j}$ are parameters. Popular candidates for $\Theta$ include monomial bases, rational functions, and Fourier bases. There exist several methods to identify nonzero coefficients $p_{j}^{i}\neq 0$, thereby learning which functions $\theta_{j}(\mathbf{x})$ influence the dynamics of state $x_{i}$, including Bayesian variable selection [7], best subset selection [8, 9], and sparse regression [10]. Of these, sparse regression has gained popularity due to its relative simplicity and low computational cost. Among the best known examples, Sparse Optimization for Nonlinear Dynamics (SINDy) finds a parsimonious model by minimizing
$L_{i}(\mathbf{p}^{i})$ $=\sum_{r=1}^{N}\|\dot{x}_{i}(t_{r})-\sum_{j=1}^{J}p_{j}^{i}\theta_{j}\left(x_{1}(t_{r}),x_{2}(t_{r}),\ldots,x_{d}(t_{r})\right)\|^{2}+\alpha\|\mathbf{p}^{i}\|_{1},\quad i=1,\ldots,d,$ (1)
@@
where $\alpha$ controls the weight of the penalty term enforcing sparsity and $N$ is the number of data samples, and $\mathbf{p}^{i}$ is a vector of parameters for the $i^{th}$ state equation. If all states $x_{i}$ are measured and the derivatives $\dot{x}_{i}$ can be measured or numerically estimated, then the model is linear with respect to the unknown parameters and hence minimization can be performed through alternating least-squares and hard-thresholding or regularized approaches such as LASSO [11], Elastic-net [12], or SR3 [13]. Prior work has addressed the challenge of applying SINDy on noisy data, including smoothing splines [14], interpolation [15], and integral formulations of sparse regression (e.g., weak SINDy [16, 17], D-SINDy [18], etc.).
@@
Such sparse regression methods are effective for applications in which the differential variables are identified, all variables are measured, and the model is linear in the unknown parameters $\mathbf{p}$. However, for complex systems described by DAEs, one or more of these assumptions break down. Recent approaches have used nonlinear optimization for sparse-model selection in systems with partial measurements [19, 20, 21, 22]. Yet, they rely on the assumption that the system is well-described by ordinary differential equations (ODEs) and do not explicitly incorporate algebraic constraints. Other groups have focused on model discovery for DAE systems but only when the conservation laws, network topology, or other algebraic constraints are known [23, 24, 25]. Some DAEs can be converted to ODEs through substitution of the algebraic equations, reducing the dimension of state space and producing ODEs with rational functions on the right-hand side. Implicit-SINDy [26], and the more computationally efficient parallel implicit SINDy (PI-SINDy) [27], have demonstrated sparse-model selection for ODEs with rational functions. However, these methods require that the differential variables are identified a priori (see Section 2.2 for further discussion). A related concept was introduced in [28], where principal component analysis (PCA) is used to eliminate correlated features from the library by retaining only those with the highest loadings in the first few principal components. However, this approach does not necessarily ensure a compact library, as the principal components may have uniformly distributed loadings unless a sparse PCA method is applied. No existing methods address the case where all states are measured, but the identity of the differential and algebraic variables is unknown, and the constraints must be identified.
@@
In this paper, we propose Sparse Optimization for Differential-Algebraic Systems (SODAs), a novel approach for discovering DAE systems by sequentially identifying the algebraic equations and the differential equations. First, we iteratively identify algebraic relationships while partially parallelizing the optimization. Once these relationships are discovered, we refine the candidate function library to mitigate multicollinearity induced by algebraic relationships. Next, the dynamic component of the DAE system is discovered using the refined candidate function library and a selection of existing methods for ODE discovery ([10, 16, 29, 30]). This sequential approach improves the condition number of the library by eliminating redundant terms and restricts the challenge posed by noisy numerical derivatives in the ODE discovery phase. We employ a singular value decomposition (SVD) analysis of the candidate library as a guiding tool to determine the stopping criteria for the refinement process under high-noise conditions. The performance of our method is illustrated using numerous DAE systems, including chemical reaction networks (CRNs), multi-body mechanical systems, and power grids.
@@
The paper is organized as follows. Section 2 provides a brief introduction to DAEs and the challenges associated with the existing SINDy framework for their discovery. The SODAs method is presented in Section 3, together with a detailed algorithm, a Python implementation of SODAs (called DaeFinder), and an example application to a simple CRN. In Section 4, we demonstrate: 1) successful rediscovery of the DAEs used to simulate time series of chemical concentrations in CRNs, 2) discovery of a reduced coordinate system and subsequent discovery of ODEs from pixel data obtained from real-time footage of experiments and video recordings of animated single and double pendulums, and 3) rediscovery of the DAEs used in power grid modeling for 3 IEEE benchmark cases. Finally, in Section 5, we address various challenges associated with our method and discuss possible improvements.
@@ 2.1 Differential-Algebraic Equations
In the types of physical systems in which we are interested, DAEs commonly arise from quasi-steady state approximations, conservation laws, and physical constraints, as seen in the modeling of chemical reactions, power grids, and multi-body systems [31]. When DAEs arise from quasi-steady state approximations [32, 33], where fast variables are assumed to reach equilibrium instantaneously relative to other variables operating in slower dynamics, it is often challenging to identify which variables are in quasi-steady state and are, therefore, algebraic rather than differential.
@@
where $\mathbf{F}=(f_{1},\ldots,f_{m})$ and $\mathbf{G}=(g_{1},\ldots,g_{d-m})$ represent the differential and algebraic parts of the system, respectively. Note that time derivatives, up to order $c$, appear only in the differential state variables $\mathbf{y}$, while the algebraic variables $\mathbf{z}$ are continuous on $[0,T]$. DAEs can also be expressed in an implicit form, where the algebraic constraints are not separated, and our algorithm is designed to work for both forms. Note that the implicit form can be converted into a semi-explicit form and also that the latter form is often more natural for many applications, as it reflects the physical significance of the explicit algebraic constraints present in these contexts. We introduce our method using the semi-explicit form.
The DAE system (2)–(3) is well-posed depending on certain regularity conditions on $\mathbf{F}$ and $\mathbf{G}$ as well as the existence of compatible initial conditions $\mathbf{x}_{0}$ satisfying the algebraic constraints (3) (see [34], Chapter 2). DAEs can be further classified based on indices related to the easiness of conversion to an ODE system (via subsequent differentiation) or based on the complexity of the numerical solution. For a detailed analysis of the solvability of nonlinear DAE systems and their classification based on various indices, see [34] and [35]. We assume that the DAE system to be discovered from data is well-posed and that the state-space trajectories $\mathbf{x}(t)$ are smooth enough on the interval $[0,T]$.
@@
$\frac{d[S]}{dt}=-k_{1}[E][S]+k_{-1}[ES],$ (4)
$\frac{d[P]}{dt}=k_{\text{cat}}[ES],$ (5)
$[E]+[ES]-E^{\text{tot}}=0,$ (6)
$\frac{d[ES]}{dt}=(k_{-1}+k_{\text{cat}})[ES]-k_{1}[E][S]\approx 0,$ (7)
where $[S],[E],[ES],[P]$ are concentrations of substrate, enzyme, enzyme-substrate complex, and products of a CRN, respectively. The enzyme-substrate complex $[ES]$ can be assumed to be in quasi-steady state under many experimental conditions [36]. Once (7) is assumed to be in quasi-steady state, the differential variables are $[S]$ and $[P]$, and the algebraic variables are $[E]$ and $[ES]$.
@@ 2.2 Data-driven discovery of equivalent ODE systems
This reduction is performed regularly in many fields, including chemical, mechanical, and electrical engineering [31, 32], to reformulate models in a manner compatible with well-established techniques for ODEs, such as numerical solvers and control methods. The reduction involves eliminating state variables using algebraic constraints and can lead to more complicated expressions in (8) as compared to (2)–(3). For example, the system (4)–(7) can be reduced using (6) and (7) to eliminate $[E]$ and $[ES]$ resulting in
$\frac{d[S]}{dt}$ $=-\frac{k_{1}k_{\text{cat}}E^{\text{tot}}[S]}{k_{-1}+k_{\text{cat}}+k_{1}[S]}.$ (9)
These are the canonical Michealis-Menten kinetics [37, 38]. More generally, DAE systems of monomial terms will generally reduce to the form
@@
Although system (10) is now converted to an ODE form, the optimization cost (1) formulated in SINDy is not a convex problem anymore because the functions are rational and hence the model is not linear in the coefficients p. This nonlinear optimization can be challenging due to various factors including non-convexity, poles of rational functions, and stability issues for techniques that use an ODE integration as part of the loss function evaluation. To address this issue, implicit-SINDy  [26] reformulated the reduced ODEs of the form (10) as an implicit ODE and then applied sparse optimization. Expressed implicitly, (10) becomes
@@
The PI-SINDy method [27] improves upon the implicit-SINDy method by performing a partial combinatorial search: each library element in $\tilde{\Theta}$ is moved to the right-hand side of the ODE and standard sparse regression is applied assuming the moved term is part of the true equation. Solving the $2J$ convex, sparse linear regressions can also be parallelized in this approach. Notwithstanding, these implicit methods still require large amounts of data, and the robustness to noise is often insufficient for many practical applications. Specifically, PI-SINDy was applied to the Michaelis-Menten kinetics model using a large dataset consisting of 150,000 data points for training and testing, generated from 3,000 randomly sampled initial conditions. Despite this large volume of data, the method was shown to handle only Gaussian noise up to a magnitude of $\sigma=0.04$ (Figure 3 in [27]), which corresponds to a noise level of less than 1% given that the initial condition for the chemical concentration ranged between 0 and 12.5 in magnitude.

### v2

@@ Abstract
Differential-algebraic equations (DAEs) integrate ordinary differential equations (ODEs) with algebraic constraints, providing a fundamental framework for developing models of dynamical systems characterized by timescale separation, conservation laws, and physical constraints. While sparse optimization has revolutionized model development by allowing data-driven discovery of parsimonious models from a library of possible equations, existing approaches for dynamical systems assume DAEs can be reduced to ODEs by eliminating variables before model discovery. This assumption limits the applicability of such methods for DAE systems with unknown constraints and time scales. We introduce Sparse Optimization for Differential-Algebraic Systems (SODAs), a data-driven method for the identification of DAEs in their explicit form. By discovering the algebraic and dynamic components sequentially without prior identification of the algebraic variables, this approach leads to a sequence of convex optimization problems. It has the advantage of discovering interpretable models that preserve the structure of the underlying physical system. To this end, SODAs improves numerical stability when handling high correlations between library terms, caused by near-perfect algebraic relationships, by iteratively refining the conditioning of the candidate library. We demonstrate the performance of our method on biological, mechanical, and electrical systems, showcasing its robustness to noise in both simulated time series and real-time experimental data.
@@ 1 Introduction
Data-driven discovery of dynamical models of complex systems promises to rapidly accelerate our understanding, from learning new equations governing cellular biomechanics [Schmitt et al. [65]] to the automatic generation of control models for robotics [Amini et al. [2], Taylor et al. [74]]. A major challenge of many complex dynamical systems is the separation of timescales and the integration of physical constraints and conservation laws. Timescale separation is ubiquitous in complex biological and physical systems, from metabolic networks to multi-body robotics and power grids. Even in well-studied systems such as those following mass-action kinetics and deterministic Newtonian physics, numerical simulation and parameter estimation are challenging due to the inherent stiffness of such systems [Higham and Trefethen [30], Shampine and Gear [68], Byrne and Hindmarsh [18]]. In this context, model selection is an even greater challenge as the structural form of the model, including its conservation laws and relevant timescales, may be unknown. In this work, we propose a method for data-driven discovery of complex systems by designing sparse-model selection within a differential-algebraic equation (DAE) framework, where conservation laws and quasi-steady-state equations are learned from time-series data as algebraic equations.
@@
Model identification or selection focuses on the general challenge of discovering the functional form of a model using data. For many dynamical systems, including those following mass-action kinetics or electromechanical conservation laws, it is reasonable to enumerate a library of possible interaction terms or candidate functions $\Theta=\{\theta_{1}(\mathbf{x}),\ldots,\theta_{J}(\mathbf{x})\}$, each representing a meaningful physical quantity as a function of the system state $\mathbf{x}(t)$. The key challenge is to downselect the most suitable subset of terms from this library. For example, a system of differential equations can often be described using a linear combination of nonlinear functions: $\dot{x}_{i}=\sum_{j=1}^{J}p_{j}^{i}\theta_{j}(\mathbf{x}),$ for each state variable $i=1,\ldots,d$, where $p_{i}^{j}$ are parameters. Popular candidates for $\Theta$ include monomial bases, rational functions, and Fourier bases. There exist several methods to identify nonzero coefficients $p_{j}^{i}\neq 0$, thereby learning which functions $\theta_{j}(\mathbf{x})$ influence the dynamics of state $x_{i}$, including Bayesian variable selection [Mitchell and Beauchamp [50]], best subset selection [Miller [48], Bertsimas et al. [8]], and sparse regression [Brunton et al. [17]]. Of these, sparse regression has gained popularity due to its relative simplicity and low computational cost. Among the best known examples, Sparse Optimization for Nonlinear Dynamics (SINDy) finds a parsimonious model by minimizing
$L_{i}(\mathbf{p}^{i})$ $=\sum_{r=1}^{N}\|\dot{x}_{i}(t_{r})-\sum_{j=1}^{J}p_{j}^{i}\theta_{j}\left(x_{1}(t_{r}),x_{2}(t_{r}),\ldots,x_{d}(t_{r})\right)\|^{2}+\alpha\|\mathbf{p}^{i}\|_{0},\quad i=1,\ldots,d,$ (1)
@@
where $\alpha$ controls the weight of the penalty term enforcing sparsity and $N$ is the number of data samples, and $\mathbf{p}^{i}$ is a vector of parameters for the $i^{th}$ state equation. If all states $x_{i}$ are measured and the derivatives $\dot{x}_{i}$ can be measured or numerically estimated, then the model is linear with respect to the unknown parameters and hence minimization can be performed through alternating least-squares and hard-thresholding or regularized approaches such as LASSO [Tibshirani [75]], Elastic-net [Zou and Hastie [87]], or SR3 [Zheng et al. [85]]. Some of these methods replace the $l_{0}$ constraint in (1) with either an $l_{1}$ or $l_{2}$ penalty, which leads to a convex optimization problem and is often combined with sequential thresholding. Each of these methods has its own advantages and disadvantages. Previous work, [Pathak and Ma [55]], has shown the sub-optimality of LASSO compared to the soft-thresholded ordinary least squares estimator (STOLS). Moreover, [Zheng et al. [86]] shows that hard thresholding combined with $l_{2}$ shrinkage can yield strong oracle-type properties and, in certain settings (particularly orthonormal designs), behaves similarly to an $l_{0}$ penalty. Model selection methods such as the one discussed in [He et al. [29]] attempt to solve the non-relaxed, non-convex $l_{0}$ optimization problem using a subspace pursuit approach, though theoretical guarantees for such methods require assumptions that are often not met in practice. Prior work has addressed the challenge of applying SINDy on noisy data, including smoothing splines [Sun et al. [73]], interpolation [Van Breugel et al. [77]], and integral formulations of sparse regression (e.g., weak SINDy [Messenger and Bortz [46, 44]], D-SINDy [Wentz and Doostan [80]], etc.).
@@
Such sparse regression methods are effective for applications in which the differential variables are identified, all variables are measured, and the model is linear in the unknown parameters $\mathbf{p}$. However, for complex systems described by DAEs, one or more of these assumptions break down. Recent approaches have used nonlinear optimization for sparse-model selection in systems with partial measurements [Ribera et al. [60], Bakarji et al. [5], Stepaniants et al. [72], Lu et al. [39]]. Yet, they rely on the assumption that the system is well-described by ordinary differential equations (ODEs) and do not explicitly incorporate algebraic constraints. Other groups have focused on model discovery for DAE systems but only when the conservation laws, network topology, or other algebraic constraints are known [Nandakumar et al. [52, 53], Sarić et al. [64]]. Some DAEs can be converted to ODEs through substitution of the algebraic equations, reducing the dimension of state space and producing ODEs with rational functions on the right-hand side. Implicit-SINDy [Mangan et al. [42]], and the more computationally efficient parallel implicit SINDy (SINDy-PI) [Kaheman et al. [33]], have demonstrated sparse-model selection for ODEs with rational functions. However, these methods require that the differential variables are identified a priori (see Section 2.2 for further discussion). A related concept was introduced in [França et al. [23]], where principal component analysis (PCA) is used to eliminate correlated features from the library by retaining only those with the highest loadings in the first few principal components. However, this approach does not necessarily ensure a compact library, as the principal components may have uniformly distributed loadings unless a sparse PCA method is applied. No existing methods address the case where all states are measured, but the identity of the differential and algebraic variables is unknown, and the constraints must be identified.
@@
In this paper, we propose Sparse Optimization for Differential-Algebraic Systems (SODAs), a novel approach for discovering DAE systems by sequentially identifying the algebraic equations and the differential equations. First, we iteratively identify algebraic relationships while partially parallelizing the optimization. Once these relationships are discovered, we refine the candidate function library to mitigate multicollinearity induced by algebraic relationships. Next, the dynamic component of the DAE system is discovered using the refined candidate function library and a selection of existing methods for ODE discovery ([Brunton et al. [17], Messenger and Bortz [46], Kang et al. [34], He et al. [29]]). This sequential approach improves the condition number of the library by eliminating redundant terms and restricts the challenge posed by noisy numerical derivatives in the ODE discovery phase. We employ a singular value decomposition (SVD) analysis of the candidate library as a guiding tool to determine the stopping criteria for the refinement process under high-noise conditions. The performance of our method is illustrated using numerous DAE systems, including chemical reaction networks (CRNs), multi-body mechanical systems, and power grids.
@@
The paper is organized as follows. Section 2 provides a brief introduction to DAEs and the challenges associated with the existing SINDy framework for their discovery. The SODAs method is presented in Section 3, together with a detailed algorithm, and an example application to a simple CRN. In Section 4, we demonstrate: 1) successful rediscovery of the DAEs used to simulate time series of chemical concentrations in CRNs, 2) rediscovery of the DAEs used in power grid modeling for 3 IEEE benchmark cases, and 3) discovery of a reduced coordinate system from pixel data obtained from real-time footage of single-pendulum experiments and from video recordings of animated single and double pendulums, which can potentially be leveraged for the subsequent discovery of ODEs. Finally, in Section 5, we address various challenges associated with our method and discuss possible improvements.
@@ 2.1 Differential-Algebraic Equations
In the types of physical systems in which we are interested, DAEs commonly arise from quasi-steady state approximations, conservation laws, and physical constraints, as seen in the modeling of chemical reactions, power grids, and multi-body systems [Campbell et al. [19]]. When DAEs arise from quasi-steady state approximations [Turanyi et al. [76], Bothe and Pierre [11]], where fast variables are assumed to reach equilibrium instantaneously relative to other variables operating in slower dynamics, it is often challenging to identify which variables are in quasi-steady state and are, therefore, algebraic rather than differential.
@@
where $\mathbf{F}=(f_{1},\ldots,f_{m})$ and $\mathbf{G}=(g_{1},\ldots,g_{d-m})$ represent the differential and algebraic parts of the system, respectively. Note that time derivatives, up to order $c$, appear only in the differential state variables $\mathbf{y}$, while the algebraic variables $\mathbf{z}$ are continuous on $[0,T]$. DAEs can also be expressed in an implicit form, where the algebraic constraints are not separated, and our algorithm is designed to work for both forms. Note that the implicit form can be converted into a semi-explicit form and also that the latter form is often more natural for many applications, as it reflects the physical significance of the explicit algebraic constraints present in these contexts. All DAE examples discussed here and used to test the algorithm have a semi-explicit form.
The DAE system (2)–(3) is well-posed depending on certain regularity conditions on $\mathbf{F}$ and $\mathbf{G}$ as well as the existence of compatible initial conditions $\mathbf{x}_{0}$ satisfying the algebraic constraints (3) (see [Brenan et al. [14]], Chapter 2). DAEs can be further classified based on indices related to the easiness of conversion to an ODE system (via subsequent differentiation) or based on the complexity of the numerical solution. For a detailed analysis of the solvability of nonlinear DAE systems and their classification based on various indices, see [Brenan et al. [14]] and [Rabier and Rheinboldt [59]]. We assume that the DAE system to be discovered from data is well-posed and that the state-space trajectories $\mathbf{x}(t)$ are smooth enough on the interval $[0,T]$.
@@
$\frac{d[A]}{dt}=-k_{1}[E_{1}][A]+k_{2}[AE_{1}],$ (4)
$\frac{d[B]}{dt}=k_{3}[AE_{1}],$ (5)
$[E_{1}]+[AE_{1}]-E_{1}^{\text{tot}}=0,$ (6)
$\frac{d[AE_{1}]}{dt}=-(k_{2}+k_{3})[ES]+k_{1}[E][S]\approx 0,$ (7)
where $A],[E_{1}],[AE_{1}],[B]$ are concentrations of substrate, enzyme, enzyme-substrate complex, and products of a CRN, respectively. The enzyme-substrate complex $AE_{1}]$ can be assumed to be in quasi-steady state under many experimental conditions [Segel and Slemrod [66]]. Once (7) is assumed to be in quasi-steady state, the differential variables are $[A]$ and $[B]$, and the algebraic variables are $E_{1}]$ and $AE_{1}]$.
@@ 2.2 Data-driven discovery of equivalent ODE systems
This reduction is performed regularly in many fields, including chemical, mechanical, and electrical engineering [Campbell et al. [19], Turanyi et al. [76]], to reformulate models in a manner compatible with well-established techniques for ODEs, such as numerical solvers and control methods. The reduction involves eliminating state variables using algebraic constraints and can lead to more complicated expressions in (8) as compared to (2)–(3). For example, the system (4)–(7) can be reduced using (6) and (7) to eliminate $E_{1}]$ and $AE_{1}]$ resulting in
$\frac{d[A]}{dt}$ $=-\frac{k_{1}k_{3}E_{1}^{\text{tot}}[A]}{k_{2}+k_{3}+k_{1}[A]}.$ (9)
These are the canonical Michealis-Menten kinetics [Johnson and Goody [32], Srinivasan [71]]. More generally, DAE systems of monomial terms can often be reduced to the following form, especially when the algebraic variables can be explicitly solved from the algebraic equations
@@
Although system (10) is now converted to an ODE form, the optimization cost (1) formulated in SINDy is not a convex problem anymore because the functions are rational and hence the model is not linear in the coefficients p. This nonlinear optimization can be challenging due to various factors including non-convexity, poles of rational functions, and stability issues for techniques that use an ODE integration as part of the loss function evaluation. To address this issue, implicit-SINDy [Mangan et al. [42]] reformulated the reduced ODEs of the form (10) as an implicit ODE and then applied sparse optimization. Expressed implicitly, (10) becomes
@@
The SINDy-PI method [Kaheman et al. [33]] improves upon the implicit-SINDy method by performing a partial combinatorial search: each library element in $\tilde{\Theta}$ is moved to the right-hand side of the ODE and standard sparse regression is applied assuming the moved term is part of the true equation. Solving the $2J$ convex, sparse linear regressions can also be parallelized in this approach. Notwithstanding, these implicit methods still require large amounts of data, and the robustness to noise is often insufficient for many practical applications. Specifically, SINDy-PI was applied to the Michaelis-Menten kinetics model using a large dataset consisting of 150,000 data points for training and testing, generated from 3,000 randomly sampled initial conditions. Despite this large volume of data, the method was shown to handle only Gaussian noise up to a magnitude of $\sigma=0.04$ (Figure 3 in [Kaheman et al. [33]]), which corresponds to a noise level of less than 1% given that the initial condition for the chemical concentration ranged between 0 and 12.5 in magnitude.